# From Ground to Sky: Architectures, Applications, and Challenges Shaping Low-Altitude Wireless Networks

### v2

@@ IV-B1 Simulations
We consider a simulation of the studied case with a fleet of five drones in the LAWN. The drones’ initial positions are drawn uniformly inside a $1500\times 1500\times 40\;\text{m}^{3}$ volume and they are cruising at $[10,15]$m/s. The desired formation is a $40$m-radius pentagon. The control plane runs a $100$Hz leader–follower model predictive controller synchronized by micro-second beacons over a sub-GHz C2 link. For payload traffic, the data plane employs a DD waveform. The transmitted DD waveform is reflected by obstacles and ISAC echoes are adopted to sense the environment. the detected obstacles will be fed to the model predictive controller. Fig. 4 plots the bit-error-rate (BER) for $1,000$ independent Monte Carlo simulations. The DD waveform significantly outperforms OFDM. In addition, swarm converges from random initial positions to the desired formation within $15$s on average.
@@ IV-B2 Real-world Experiment
We also conduct a real-world experiment with four DJI Tello drones in a $10$×$10$m motion-capture arena. A trolley-mounted laptop serves as the ground LAWN node. From arbitrary starts, the drones form a square formation with four translational motions, namely up, forward, down, and backward. The laptop commands these manoeuvres to the designated leader. Then, the three followers maintain fixed displacement offsets, allowing the square to translate rigidly while holding formation throughout the motion.
@@ V Cross-layer Challenges of the LAWN
In previous sections, we have discussed the four‑plane architecture, enabling technologies, and potential applications of LAWNs. Nevertheless, a real-world LAWN must cope with technical limits, aviation safety rules, and policy constraints.
@@
V-B Control with Low Latency in Congested Skies
With the rapid development of LAWN, it is expected that the low-altitude airspace will become increasingly crowded. In such scenarios, drones may fly in close formations, maneuver rapidly to avoid obstacles, raising the demand of end-to-end control loops with extremely stringent latency constraints. On the one hand, short frame structures, minimal preambles, and Doppler-resilient waveforms will be adopted for reducing transmission delays. On the other hand, the data plane scheduler should swiftly re-allocate resources for critical C2 message and non-urgent data streams. Developing congestion control algorithms and adaptive network slicing schemes that are capable of dynamically adjusting priorities and resources within microseconds is of great importance.
V-C Cyber-Physical Security
The close integration of cyber and physical components in LAWN poses critical security challenges. Drones in low-altitude airspace are especially vulnerable to attacks such as GNSS spoofing, C2 injection, and communication jamming, which may result in physical threats and impact public safety. To ensure cyber-physical security in LAWN, developing robust and computationally efficient cryptographic protocols is necessary. Moreover, designing onboard-embedded AI algorithms to detect abnormal or malicious behavior can prevent LAWN from severe incidents. In addition, LAWN can harness physical layer security methods, such as artificial-noise beamforming, cooperative jamming, and adaptive waveform design, to prevent the eavesdropping and jamming directly at the radio interface with minimal computational overhead.
V-D Scalability Versus Battery Limits
@@
V-E Traffic Management and Global Regulations
@@ VI Conclusion and Outlook
In this article, we have discussed the concept of LAWN, which offers a new framework for realizing intelligent, resilient, and multifunctional wireless systems in the low-altitude airspace. By tightly integrating data transmission, control, environmental sensing, and computing, the LAWN extends the traditional role of drones as communication relays or remote sensors. Instead, it establishes a reconfigurable, three-dimensional infrastructure capable of adapting dynamically to mission objectives, environmental conditions, and airspace constraints. Realizing this vision requires cross-disciplinary collaboration across wireless communications, control theory, robotics, embedded AI, and aviation regulation. Looking ahead, we anticipate the LAWN to play a foundational role in the emerging low-altitude ecosystems. It has the potential to transform how cities manage logistics, how rural areas access services, and how societies monitor and interact with the environment. This article aims to provide both a conceptual foundation and a research roadmap for advancing LAWN, encouraging the academic and industrial communities to jointly shape the future of autonomous aerial networks.

### v3

@@ IV-B1 Simulations
We consider a simulation of the studied case with a fleet of five drones in the LAWN. The drones’ initial positions are drawn uniformly inside a $1500\times 1500\times 40\;\text{m}^{3}$ volume and they are cruising at [10,15][10,15]m/s. The desired formation is a 4040m-radius pentagon. The control plane runs a 100100Hz leader–follower model predictive controller synchronized by micro-second beacons over a sub-GHz C2 link. For payload traffic, the data plane employs a DD waveform. The transmitted DD waveform is reflected by obstacles and ISAC echoes are adopted to sense the environment. the detected obstacles will be fed to the model predictive controller. Fig. 4 plots the bit-error-rate (BER) for $1,000$ independent Monte Carlo simulations. The DD waveform significantly outperforms OFDM. In addition, swarm converges from random initial positions to the desired formation within $15$s on average.
@@ IV-B2 Real-world Experiment
We also conduct a real-world experiment with four DJI Tello drones in a 1010×1010m motion-capture arena. A trolley-mounted laptop serves as the ground LAWN node. From arbitrary starts, the drones form a square formation with four translational motions, namely up, forward, down, and backward. The laptop commands these manoeuvres to the designated leader. Then, the three followers maintain fixed displacement offsets, allowing the square to translate rigidly while holding formation throughout the motion.
TABLE II: Roadmap of Cross-Layer Challenges for LAWNs
Challenge Potential solutions Performance metrics
3D spectrum coexistence: Altitude-/orientation-dependent interference and fast beam re-pointing under mixed frequency bands yield rapidly varying interference footprints Dynamic 3D interference modeling, predictive beam/power scheduling, sensing-assisted spectrum awareness SINR, outage, handover interruption
Time synchronization & localization integrity: GNSS degradation, multipath, and clock drift lead to biased locations and timings in drone swarms Network-assisted synchronization, integrity monitoring, cooperative localization Synchronization error, integrity-bounded position error/undetected failure, spoofing detection time
Low-latency control in congested skies: Dense formations and rapid maneuvers demand extreme tail-latency and ultra-high reliability C2 scheduling/isolation, packet duplication + multi-connectivity, fast congestion control Tail latency, C2 freshness (age of information), reliability under blockage
Cyber-physical security: GNSS spoofing, C2 injection, and jamming can trigger immediate physical hazards; mesh routing and edge offloading expand attack surface under tight compute/energy budgets Lightweight authentication, physical layer security, anomaly detection Detection delay, false alarm&miss, time-to-safe-state
Scalability vs. battery limits: More links, sensing, and inference increase energy consumption, and relay imbalance can break mesh connectivity under limited battery capacity Energy-aware scheduling/routing, relay-load balancing, trajectory planning, opportunistic charging coordination Mission time, energy per control update, supported density.
Traffic management & regulations:UTM must scale to support massive autonomous operations, while regional policy constraints directly shape network topology and service provisioning Hybrid UTM, standardized UTM–LAWN interfaces, policy-aware routing Compliance rate, conflict resolution time
Explainable & robust AI: Black-box decision making under distribution shifts or adversarial inputs undermine trust, while errors can rapidly propagate through multi-agent closed-loop interactions Uncertainty-aware inference/planning, XAI summaries/attribution, adversarial robustness and credibility checks Calibration, worst-case violations, AI-induced failure rate
Intermittent connectivity & partitioned swarms: Blockage and mobility can fragment the aerial mesh Opportunistic relaying, partition-aware task allocation, redundant safety signaling with local fallback control Partition frequency&duration, service continuity, mission utility.
@@ V Cross-layer Challenges of the LAWN
In previous sections, we have discussed the four‑plane architecture, enabling technologies, and potential applications of LAWNs. In practice, a LAWN must allocate spectrum, computation, and energy under uncertainty while preserving mission-grade reliability and regulatory compliance. Below, we outline key challenges that can serve as a roadmap for future research and standardization.
@@
V-B Time Synchronization and Localization Integrity
Accurate time synchronization and reliable localization are foundational to many LAWN function. In practice, however, low-altitude operations frequently take place in GNSS-challenging environments such as urban canyons and post-disaster areas with degraded infrastructure. Under these conditions, clock drift, multipath, and intermittent visibility of satellites can lead to biased position estimates and inconsistent timing across drones. Such errors may propagate through the sensing and control planes, degrading collision-avoidance performance. Moreover, malicious threats such as spoofing can further degrade timing and positioning, making integrity monitoring as important as raw accuracy. Addressing this challenge requires multi-sensor fusion with network-assisted synchronization over D2D/D2G links, together with integrity measures that can detect and bound localization failures in real time.
V-C Control with Low Latency in Congested Skies
With the rapid development of LAWN, it is expected that the low-altitude airspace will become increasingly crowded. In such scenarios, drones may fly in close formations, maneuver rapidly to avoid obstacles, raising the demand of end-to-end control loops with extremely stringent latency constraints[28]. On the one hand, short frame structures, minimal preambles, and Doppler-resilient waveforms will be adopted for reducing transmission delays. On the other hand, the data plane scheduler should swiftly re-allocate resources for critical C2 message and non-urgent data streams. Developing congestion control algorithms and adaptive network slicing schemes that are capable of dynamically adjusting priorities and resources within microseconds is of great importance.
V-D Cyber-Physical Security
The close integration of cyber and physical components in LAWN poses critical security challenges. Drones in low-altitude airspace are especially vulnerable to attacks such as GNSS spoofing, C2 injection, and communication jamming, which may result in physical threats and impact public safety. To ensure cyber-physical security in LAWN, developing robust and computationally efficient cryptographic protocols is necessary. Moreover, designing onboard-embedded AI algorithms to detect abnormal or malicious behavior can prevent LAWN from severe incidents. In addition, LAWN can harness physical layer security methods, such as artificial-noise beamforming, cooperative jamming, adaptive waveform design and covert communications, to prevent the eavesdropping and jamming directly at the radio interface with minimal computational overhead.
V-E Scalability Versus Battery Limits
@@
V-F Resilience to Intermittent Connectivity and Partitioned Swarms
A defining feature of LAWNs is rapid topology change under mobility and blockage. In cluttered low-altitude environments, LoS links can be abruptly blocked and the aerial mesh may temporarily split into multiple partitions. These intermittent and partitioned conditions challenge conventional routing and scheduling, which often assume persistent end-to-end paths and stable neighbor relations. Robust LAWN operation therefore requires connectivity-aware and disruption-tolerant mechanisms that can gracefully degrade service, such as opportunistic relaying and partition-aware task allocation that avoids assigning tightly coupled tasks to disconnected sub-swarms.
V-G Traffic Management and Global Regulations
@@ VI Conclusion and Outlook
In this article, we have discussed the concept of LAWN, which offers a new framework for realizing intelligent, resilient, and multifunctional wireless systems in the low-altitude airspace. By tightly integrating data transmission, control, environmental sensing, and intelligence/computing, the LAWN extends the traditional role of drones as communication relays or remote sensors. Instead, it establishes a reconfigurable, three-dimensional infrastructure capable of adapting dynamically to mission objectives, environmental conditions, and airspace constraints. Realizing this vision requires cross-disciplinary collaboration across wireless communications, control theory, robotics, embedded AI, and aviation regulation. Looking ahead, we anticipate the LAWN to play a foundational role in the emerging low-altitude ecosystems. It has the potential to transform how cities manage logistics, how rural areas access services, and how societies monitor and interact with the environment. This article aims to provide both a conceptual foundation and a research roadmap for advancing LAWN, encouraging the academic and industrial communities to jointly shape the future of autonomous aerial networks.